# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ 2.1 Framework and first results
Let $\sigma=(f,g)$ be a sequence of surfaces of the same degree, say $d_{f}=d_{g}=p$, called a pencil of surfaces. It defines a 11-parameter family of projective surfaces as:
@@
and each hypersurface $V(z_{0}f+z_{1}g)$ is called a member of the pencil $V_{\sigma}$. The set where each surface is singular is the vanishing locus of a polynomial $p=p(z)$ (the discriminant) inside $\mathbb{P}^{1}$, so it must be either a finite subset or the whole family. When this set is finite, the pencil is called a regular pencil, otherwise is called irregular.
In [undefj, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}$ for pencils of surfaces as the union of singular loci:
@@
Hence, whenever a normal sequence $\sigma=(f,g)$ is a regular pencil where every singular member has isolated singularities, we conclude $\dim(\Xi_{\sigma})=0$ from the description above, and thus $m(\sigma)=0$.
@@
$\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
given by the Buchsbaum–Rim complex.
@@ 2.2 Plane curves as pairs of surfaces
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$. Denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme, given as the zero locus of the Jacobian ideal of $g$.
The matrix $\nabla\sigma$ will be given by
@@
the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
@@
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$. Denoting $i:H\hookrightarrow\mathbb{P}^{2}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
@@
where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:H\hookrightarrow\mathbb{P}^{3}$ and thus
@@
$=1-d[H]+(d^{2}-\tau(X))[H]^{2},$
and since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
Moreover, doing the analogous comparison for the total classes of the twist ${\mathcal{T}}_{\sigma}(e)$, we obtain
$i^{*}c({\mathcal{T}}_{\sigma}(e))=1+(e-d)[H]+\operatorname{Bour}(\sigma)[H]^{2}=1+(e-d)[H]+\operatorname{Bour}(X)[H]^{2}=c(i^{*}{\mathcal{T}}_{\sigma}(e)),$
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$. Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
A more algebraic way to see this is through the associativity formula. Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$, with the effect of adding a variable to the graded ring of the plane $\mathbb{P}^{2}\hookrightarrow\mathbb{P}^{3}$, and the associated map of modules is given by the quotient map
$0\to(x_{3})\to\kappa[x_{0},x_{1},x_{2},x_{3}]\to\kappa[x_{0},x_{1},x_{2}]\to 0.$
Then, for any associated prime $P$ of the Jacobian ideal of $X$ at $\kappa[x_{0},x_{1},x_{2}]$, the extension will satisfy
$\dim R/(R\cdot P)$ $=\dim(\kappa[x_{0},x_{1},x_{2}]/P)+1$
$\deg R/(R\cdot P)$ $=\deg(\kappa[x_{0},x_{1},x_{2}]/P).$
since the generators do not include $x_{3}$. Moreover, the length of the modules over $P$ will be preserved, as the element $x_{3}\not\in P$ will be invertible. Together with the associativity formula, we obtain:
$\tau(X)$ $=\sum_{P\in\operatorname{Ann}(\mathcal{O}_{\operatorname{Sing}X}(d)),\dim(\kappa[x_{0},x_{1},x_{2}]/P)=1}\operatorname{{length}}(\mathcal{O}_{\operatorname{Sing}X}(d))_{P}\cdot\deg(\kappa[x_{0},x_{1},x_{2}]/P)$
$=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(\kappa[x_{0},x_{1},x_{2},x_{3}]/P)=2}\operatorname{{length}}(\mathcal{Q}_{\sigma})_{P}\cdot\deg(R/P)=m(\sigma).$
The following example of the nature above is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ but $m(\sigma)\neq 0$, showing that a converse for 8, $(a)$ is not possible without assuming the Buchsbaum–Rim resolution.
@@
2.3 Free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [undefr, Theorem 2.1, (c)]. The next result follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$, as it is the case for $\mathbb{P}^{2}$.
@@ 2.3 Free resolutions and the Bourbaki scheme
Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a non-zero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated to $\nu$ in a short exact sequence:
@@
    lifts for a free resolution of the form
    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0$
    for ${\mathcal{T}}_{\sigma}$.
@@
    For a minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
@@
    it induces a free resolution for $\mathcal{I}_{B}$ of the form:
@@
since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$. Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\omega(e-d)=\tilde{\omega}$. We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
@@
Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say $\sigma$ is:
@@
  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of Lemma 12, $(b)$.
Example 14 (Nearly free sequence with $d_{f}=1,d_{g}=2$).
@@
The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$. We also note that there is an easy syzygy of degree one, namely $\nu=(-x_{3},0,0,x_{0})^{T}$, hence $e=1$.
To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated to the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{P}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}_{P}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$. From this, we may rewrite the matrix as
@@
first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$. Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{P}^{2}$, and thus
@@
since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$. Thus, it follows $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$. For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
$m(\sigma)=\sum_{i=1}^{3}\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4.$
Using the Bourbaki formula with $m_{0}=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}=7$, $d=3$ and the obtained discrete invariants $m(\sigma)=4$ and $e=1$, we conclude that $\operatorname{Bour}(\sigma)=1$.
From the minimal free resolution of the ideal of a line in $\mathbb{P}^{3}$, using Lemma 12, the minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, one obtains $c_{3}({\mathcal{T}}_{\sigma})=3$.
@@
Considering the following pencil of cubics:
$\sigma=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
@@
The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals of $R$:
@@
We claim that $m(\sigma)=5$, a quantity which will be obtained using the associativity formula. For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we start by noting that the quantities:
@@
are all invertible inside the ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$. Then, we do two elementary operations:
@@
and hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
@@
is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit. Thus, we obtain $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
@@
noting that the factors $(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1}))$ and $2x_{2}x_{3}(x_{0}-x_{1})$ are units in the local ring. Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula. The main result of the next section is:
$e=1$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 2$
$e=2$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 5,$
so for $m_{0}=12$ we conclude that these two correspond to $m(\sigma)\geq 7$ and $m(\sigma)\geq 6$, respectively. Since $m(\sigma)=5$, we conclude automatically that $e\geq 3$.
Using Macaulay2, we obtain a minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the 33-syzygy kind, given by:
so $e=3$ and, with $m(\sigma)=5$, we conclude $\operatorname{Bour}(\sigma)=4$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for the Bourbaki ideal $\mathcal{I}_{B}$ will be of the form:
@@
presenting $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
@@
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then:
  • (a)
    $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
    where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
  • (b)
    if $\sigma$ is nearly free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
@@
This follows from Lemma 12, using the minimal free resolution for a line in $l\subset\mathbb{P}^{3}$ as the intersection of two planes:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for a choice of syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. The isolated zeros of the minors will be in the support of the sheaf $\operatorname{supp}({\mathcal{T}}_{\sigma})$, and these will be inside the line $l$ from the resolution obtained. ∎
@@
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [undefd], related to rational cuspidal curves. In [undefr, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [undefd] (see [undefr, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of previous works, for example [undef], [undefe] and [undefg].

### 27 Apr 2026 v4

@@ 2.1 Framework and first results
Let $\sigma=(f,g)$ with $d_{f}=d_{g}=p$. It defines a pencil of projective surfaces:
@@
where each $V(z_{0}f+z_{1}g)$ is a member of $V_{\sigma}$. The pencil is called regular when the generic member is non-singular, otherwise it is called irregular.
In [11, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ for pencils of surfaces as the union of singular loci:
@@
Hence, whenever a normal sequence $\sigma=(f,g)$ defines a regular pencil whose members have at most isolated singularities, then $\dim(\Xi_{\sigma})=0$ and $m(\sigma)=0$.
@@
$\sigma=(f,g)=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$. Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Using 9, $(a)$, it follows ${\mathcal{T}}_{\sigma}$ has a minimal free resolution given by the Buchsbaum–Rim complex:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
@@ 2.2 Plane curves as pairs of surfaces
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free homogeneous polynomial of degree $d$, with the associated curve $X=V(g)\subset\mathbb{P}^{2}=V(x_{3})$ which has at most isolated singularities, and let $\sigma=(x_{3},g)$. We consider $S=V(g)\subset\mathbb{P}^{3}$, which is the cone over the curve $X$ and denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme. The matrix $\nabla\sigma$ will be given by
@@
the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, we form the following diagram with exact columns:
@@
and by the snake lemma, there are isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$. On $\mathbb{P}^{2}$, we have the sequence defining the logarithmic tangent sheaf of the curve $X$:
$0\to{\mathcal{T}}_{g}\to\mathcal{O}_{\mathbb{P}^{2}}\xrightarrow{(\partial_{0}g,\partial_{1}g,\partial_{2}g)}\mathcal{O}_{\mathbb{P}^{3}}(d)\to 0,$
and a defined Bourbaki degree $\operatorname{Bour}(X)$ ([17]), which coincides with the Chern class $c_{2}({\mathcal{T}}_{g}(e))$ when $e=\operatorname{indeg}({\mathcal{T}}_{g})$. By the block-form of $\nabla\sigma$,
$e=\operatorname{indeg}({\mathcal{T}}_{g})=\operatorname{indeg}({\mathcal{T}}_{\sigma}).$
Denoting by $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely. Now, compare the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
@@
where $\tau(X)$ is the Tjurina number of the curve $X$. From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$, so that
@@
$=1-d[H]+(d^{2}-\tau(X))[H]^{2}.$
Since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$. Geometrically, this means that $m(\sigma)$ counts the singular lines of the cone $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does. Another perspective for this is through the associativity formula. Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$ or, in terms of rings, if $S\doteq\kappa[x_{0},x_{1},x_{2}]$, we have
$\frac{S[x_{3}]}{(\partial_{0}g,\partial_{1}g,\partial_{2}g)}\otimes_{S}\frac{S[x_{3}]}{(x_{3})}\simeq\frac{S}{(\partial_{0}g,\partial_{1}g,\partial_{2}g)},$
and then coming back to $R$ via a flat change of basis $\otimes_{\kappa}\kappa[x_{3}]$, since $\partial_{i}g\in\kappa[x_{0},x_{1},x_{2}]$. Therefore, for any associated prime $\mathfrak{p}$ of the Jacobian ideal $J_{g}\subset S$, the extension satisfies
$\dim R/(\mathfrak{p}R)$ $=\dim(S/\mathfrak{p})+1$
$\deg R/(\mathfrak{p}R)$ $=\deg(S/\mathfrak{p}),$
and each associated prime to the cone ideal $J_{g}R$ comes by an extension to an associated prime of $J_{g}$ in $S$. Now, using the associativity formula, we get:
$\tau(X)$ $=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(J_{g}),\dim(S/\mathfrak{p})=1}\operatorname{\operatorname{length}}_{S_{\mathfrak{p}}}(S/J_{g})_{\mathfrak{p}}\cdot\deg(S/\mathfrak{p})$
$=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ass}}(J_{g}R),\dim(R/\mathfrak{p})=2}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}(R/J_{g}R)_{\mathfrak{p}}\cdot\deg(R/\mathfrak{p})=m(\sigma).$
The following example is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ is maximal but $m(\sigma)\neq 0$.
@@
2.3 Free resolutions and Bourbaki schemes
In this section, after a choice of syzygy $\nu$ of minimal degree, we relate the resolutions for the ideal sheaf of its zero-locus $\mathcal{I}_{B_{\nu}}$ and for ${\mathcal{T}}_{\sigma}$. We use this lemma to characterize different classes of sequences.
@@ 2.3 Free resolutions and Bourbaki schemes
Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a nonzero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated with $\nu$ in a short exact sequence:
@@
    lifts to a free resolution
    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0.$
@@
    Every minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
@@
    induces a free resolution of $\mathcal{I}_{B}$:
@@
To summarize, item $(a)$ follows from the Horseshoe lemma since line bundles have vanishing $\operatorname{{Ext}}^{1}(-,\mathcal{O}_{\mathbb{P}^{3}})$, and $(b)$ comes from the mapping cone, after factoring the trivial factors $\mathcal{O}_{\mathbb{P}^{3}}(-e)$.
@@
since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$. Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\tilde{\omega}=\omega(e-d)$. We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
@@
Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say that $\sigma$ is:
@@
  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of 13, $(b)$.
Example 15 (Nearly-free sequence with $d_{f}=1,d_{g}=2$).
@@
The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$. We also note that $\nu=(-x_{3},0,0,x_{0})^{T}$ is a syzygy for $\nabla\sigma$ of degree one, so $e\leq 1$. We claim that $m(\sigma)=4$, which shows that $\sigma$ cannot be compressible (by Theorem A Theorem A) and hence $e=1$. Moreover, with these data, the Bourbaki degree formula yields $\operatorname{Bour}(\sigma)=1$, so this is a nearly-free sequence. To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated with the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{\mathfrak{p}}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$. From this, we may rewrite the matrix as
@@
first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$. Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{\mathfrak{p}}^{2}$, and thus
@@
since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$. Thus, it follows $\operatorname{\operatorname{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$. For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
$m(\sigma)=\sum_{i=1}^{3}\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4,$
as we wanted to show.
@@
Consider the following pencil of cubics:
$\sigma=(f,g)=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
@@
We claim that here, $m(\sigma)=5$ and that $\operatorname{Bour}(\sigma)=4\neq 1$, so $\sigma$ is not nearly free. Macaulay2 gives the following minimal free resolution for ${\mathcal{T}}_{\sigma}$:
which is of $3$-syzygy type. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals in $R$:
@@
For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we note that the elements
@@
are all invertible in the local ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$. Then, we do two elementary operations:
@@
and hence $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
@@
is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit. Thus, we obtain $\operatorname{\operatorname{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
@@
noting that the factors $(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1}))$ and $2x_{2}x_{3}(x_{0}-x_{1})$ are units in the local ring. Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula.
Using 13 with the minimal free resolution for ${\mathcal{T}}_{\sigma}$ above, after a choice of minimal degree syzygy, one gets the following resolution for $B=B_{\nu}$:
@@
which presents $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
@@
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. Moreover, when $\sigma$ is nearly free, the zero-dimensional part of $\Xi_{\sigma}$ is aligned.
@@
Note that $\operatorname{Bour}(\sigma)=1$ if and only if $B=B_{\nu}$ is a line, for a given choice of non-zero syzygy of minimal degree $\nu$. Then, the equivalence follows from 13. In general, the singular scheme of ${\mathcal{T}}_{\sigma}$ is supported at the zero-dimensional locus of $\Xi$, and this will be contained inside $B$, a fact which may be seen from the short exact sequence
$0\to\mathcal{O}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{I}_{B}(e-d)\to 0.$
@@
The notion of nearly-free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [7], related to rational cuspidal curves. In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [7] (see [17, Definition 2.17]). The notion of $3$-syzygy divisors is also present in a number of previous works, for example [1], [8] and [9].